a. Forecast the parts of the income statement and balance sheets necessary to calculate free cash flow.

Partial Income Statement for the Year Ending December 31 (Millions of Dollars)

 

Actual

Projected

Projected

Income Statement Items

2018/12/31

12/31/19

12/31/20

Net Sales

$800.0

 

 

Costs (except depreciation)

$576.0

 

 

Depreciation

$60.0

 

 

   Total operating costs

$636.0

 

 

Earning before int. & tax

$164.0

 

 

Partial Balance Sheets for December 31 (Millions of Dollars)

 

Actual

Projected

Projected

Operating Assets

2018/12/31

12/31/19

12/31/20

Cash

$8.0

 

 

Accounts receivable

$80.0

 

 

Inventories

$160.0

 

 

Net plant and equipment

$600.0

 

 

 

 

 

 

Operating Liabilities

 

 

 

Accounts Payable

$16.0

 

 

Accruals

$40.0

 

 

b.   Calculate free cash flow for each projected year. Also calculate the growth rates of free cash flow each year to ensure that there is constant growth (i.e., the same as the constant growth rate in sales) by the end of the forecast period.

 

Actual

Projected

Projected

Calculation of FCF

2018/12/31

12/31/19

12/31/20

Operating current assets

 

 

 

Operating current liabilities

 

 

 

Net operating working capital

 

 

 

Net PPE

 

 

 

Total net operating capital

 

 

 

NOPAT

 

 

 

Investment in total net operating capital

na

 

 

Free cash flow

na

 

 

Growth in FCF

na

na

 

Growth in sales

 

 

 

 

 

 

 

Income Statement for the Year Ending December 31 (Millions of Dollars) 2018 S Net Sales 800.0 Costs (except depreciation) Depreciation S 576.0 S 60.0 Total operating costs S 636.0 S Earning before int. & tax 164.0 S Less interest 32.0 Earning before taxes Taxes (40%) 132.0 52.8 S Net income before pref. div. 79.2 Preferred div. 1.4 Net income avail. for com. div. 77.9 S Common dividends 31.1 Addition to retained earnings 46.7 Number of shares (in millions) Dividends per share 10 3.11 Balance Sheets for December 31 (Millions of Dollars) Liabilities and Equity Accounts Payable Notes payable Assets 2018 2018 Cash 8.0 16.0 Short-term investments 20.0 40.0 80.0 Accounts receivable Accruals 40.0 Inventories 160.0 Total current liabilities $ 96.0 S Total current assets 268.0 Long-term bonds 300.0 Net plant and equipment 15.0 600.0 Preferred stock (Par plus PIC) S Total Assets 868.0 257.0 Retained earnings 200.0 S Common equity 457.0 Total liabilities and equit $ 868.0
Projected ratios and selected information for the current and projected years are shown below. Inputs Actual Projected Projected 2018/12/31 Projected Projected 12/31/21 12/31/19 12/31/20 12/31/22 Sales Growth Rate 15% 10% 6% 6% Costs/Sales 72% 72% 72% 72% 72% Depreciation/(Net PPE) 10% 10% 10% 10% 10% Cash/Sales 1% 1% 1% 1% 1% (Acct. Rec.)/Sales 10% 10% 10% 10% 10% Inventories/Sales 20% 20% 20% 20% 20% (Net PPE)/Sales (Acct. Pay.)/Sales 75% 75% 75% 75% 75% 2% 2% 2% 2% 2% Accruals/Sales 5% 5% 5% 5% 5% Tax rate 40% 40% 40% 40% 40% |Weighted average cost of capital (WACC) 10.5% 10.5% 10.5% 10.5% 10.5%
Answer #1

ANSWER:-

Price per share = $ 99.46

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. You can also understand the mathematics making use of [+] / [-] sign appearing in the first column. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $ mn except number of shares which is in mn and Price per share which is $ / shares.

Parameter Linkage Actual Projected Projected Projected Projected
Year 12/31/2018 12/31/2019 12/31/2020 12/31/2021 12/31/2022
Year number, n 1 2 3 4
y-o-y growth in sales gi 15% 10% 6% 6%
Net Sales A = Ai-1 x (1 + gi) 800         920.00       1,012.00       1,072.72       1,137.08
Costs B = 72% x A         662.40          728.64          772.36          818.70
Depreciation C = 10% x I (below)           69.00            75.90            80.45            85.28
Cash D = 1% x A                8.00             9.20            10.12            10.73            11.37
Acc. Receivables E = 10% x A              80.00           92.00          101.20          107.27          113.71
Inventory F = 20% x A            160.00         184.00          202.40          214.54          227.42
Accounts Payable G = 2% x A              16.00           18.40            20.24            21.45            22.74
Accruals H = 5% x A              40.00           46.00            50.60            53.64            56.85
Net PPE I = 75% x A            600.00         690.00          759.00          804.54          852.81
Working capital J = (D + E + F) - (G + H)            192.00         220.80          242.88          257.45          272.90
Increase in working capital K = Ji+1 - Ji           28.80            22.08            14.57            15.45
Capital Expenditure L = Ii + C - Ii-1         159.00          144.90          125.99          133.55
EBIT x (1 - Tax rate) M = (A - B - C) x (1 - 40%)         113.16          124.48          131.94          139.86
[+] Depreciation C           69.00            75.90            80.45            85.28
[-] Increase in working capital K           28.80            22.08            14.57            15.45
[-] Capital expenditure L         159.00          144.90          125.99          133.55
FCFF N = M + C - K - L            (5.64)            33.40            71.83            76.14
WACC R 10.5%
Terminal growth rate g 6%
Horizon value O = N2022 x (1 + g) / (R - g)       1,793.56
PV factor PVF = (1 + R)-n       0.90498        0.81898        0.74116        0.67073
PV of FCFF PV = PVF x FCFF            (5.10)            27.35            53.24            51.07
[+] PV of Horizon value PVHV = O x PVF2022       1,203.00
Value of operations (V0) V0 = PV + PVHV         1,329.56
[+] Value of short-term investments P              20.00
Total value of company Q = V0 + P         1,349.56
[-] Total value of all debt R (= Notes payables + long term bonds)            340.00
[-] Value of preferred stock S              15.00
Value of common equity T = Q - R - S            994.56
Number of shares outstanding (mn) U              10.00
Price per share ($ / share)